# Robust Methods for Multiscale Coarse Approximations of Diffusion Models in Perforated Domains

### v3

@@ 4.1 Coarse Approximation on L-Shaped Domain
In order to assess the convergence of the discrete Trefftz method, we consider two strategies regarding the refinement of the coarse partitioning. The procedure involving the reduction of the diameter of the coarse cells will be referred to as mesh refinement. The sequence of such meshes will be constructed as follows: first the background domain $D$ is partitioned into $N=(2k+1)^{2}$, $k\in\mathbb{N}$, squares, then the coarse cells $\Omega_{j}$ are generated by excluding $\Omega_{S}$. The choice of $N$ being a square of an odd number ensures the consistency of the mesh sequence in terms of the shape of the elements. The second considered refinement strategy will be referred to as edge refinement procedure, which involves subdividing the edges of an original “$3\times 3$ grid”. This edge refinement approach is illustrated by Figure 4 and is inspired by the Multiscale Hybrid-Mixed method Araya et al. (2013). We remark that with both refinement procedures, it is ensured that none of the coarse grids will have a degree of freedom located at the corner $(0,0)$. As a result, the corner singularity will be captured by the basis functions associated with the L-shaped domain. We also note that in order to improve the precision of the fine-scale finite element method, the size of the triangles is graded in the vicinity of the corner $(0,0)$. Figure 5 reports, for both mesh and edge refinement strategies, the relative error in $H^{1}$ semi-norm and $L^{2}$ norm as functions of maximal coarse edge length $H$. The black dashed line represents the typical fine-scale finite element error; here, we use $\mathbb{P}_{2}$ finite elements. The relative error and the experimental order of convergence for edge refinement procedure are equally reported in Table 1. In accordance with the error estimate (11), for the edge refinement procedure, we observe superconvergence in the energy norm with rates of approximately $3/2$ and $5/2$ for $p=1$ and $p=2$, respectively. In the $L^{2}$ norm, the observed convergence rates are approximately $3$ and $7/2$ for $p=1$ and $p=2$, respectively. In contrast to the edge refinement convergence, the convergence resulting from the mesh refinement seems to be controlled by the low global regularity of the solution. We observe the convergence rates typical for finite element methods on quasi-uniform meshes, that is, close to $2/3$ and $4/3$ in $H^{1}$ and $L^{2}$ norms, respectively.
@@ 4.3 Iterative DD and Preconditioned Krylov Methods on Small Urban Domain
We examine the performance of the two-level stationary iteration and preconditioned GMRES method over a domain based on realistic urban geometries for which the data sets were kindly provided by Métropole Nice Côte d’Azur. We focus here on a relatively small spatial frame shown on the left of Figure 1). The data frame is $160\times 160$ meters and contains two kinds of structural features representing buildings (and assimilated small elevated structures) and walls in urban data. The number of perforations associated to buildings and walls is 63 and 77, respectively. The triangulation of the domain is performed using Triangle Shewchuk (1996). This “small” model domain would take a minimum of 23  000 degrees of freedom to triangulate without imposing maximum triangle area or constraining the mesh to match the coarse partitioning. While the fine-scale mesh will change depending on $N$, the number of degrees of freedom in the mesh will be close to this number. Due to the geometric complexity of the computational domain, the mesh includes a number of very small triangles; the minimal triangle area is given as $5.83\times 10^{-13}$. We consider the model problem (1) with $f=1$, which we discretize with continuous piece-wise affine finite elements; the finite element solution is reported on the left of Figure 1.
@@
We expose on Figure 8 the convergence history of the two-level stationary iteration method with overlap $\frac{1}{20}\mathcal{H}_{j}$. We observe that the convergence of the algebraic error is robust with respect to $h$ and that the fine-scale finite element method (black lines) can be reached in a few steps. Further iterations do not improve the overall precision of the approximate solution even though the algebraic error may decrease. A stopping criterion which prevents excess iterations is warranted; such a posterioiri estimates are discussed for example in Arioli et al. (2005); Jiránek et al. (2010). As expected, the performance of the stationary iteration method considerably deteriorates (see Figure 9) in the case of minimal geometric overlap.
@@ 4.4 Scalability Tests for Preconditioned Krylov Method on Larger Urban Domain
For the sake of comparison, we also provide numerical results for the well-known Nicolaides coarse space Nicolaides (1987), made of flat-top partition of unity functions associated with the overlapping partitioning. For the Nicolaides coarse space, scalability with respect to the number of subdomains is only guaranteed if the subdomains are connected. Therefore, to generate connected subdomains, we further partition $(\Omega_{j}^{\prime})_{j}$ into a family of connected subdomains. In other words, let $m_{j}$ denote the number of disconnected components for each overlapping subdomain $\Omega_{j}^{\prime}$ and let $\Omega_{j,l}^{\prime},l=1,\ldots,m_{j}$ denote the corresponding disconnected component. Then our new overlapping partitioning contains $m=\sum_{j=1}^{N}m_{j}$ total subdomains and is given by

### v4

@@ 4.1 Coarse Approximation on L-Shaped Domain
In order to assess the convergence of the discrete Trefftz method, we consider two strategies regarding the refinement of the coarse partitioning. The procedure involving the reduction of the diameter of the coarse cells will be referred to as mesh refinement. The sequence of such meshes will be constructed as follows: first the background domain $D$ is partitioned into $N=(2k+1)^{2}$, $k\in\mathbb{N}$, squares, then the coarse cells $\Omega_{j}$ are generated by excluding $\Omega_{S}$. The choice of $N$ being a square of an odd number ensures the consistency of the mesh sequence in terms of the shape of the elements. The second considered refinement strategy will be referred to as edge refinement procedure, which involves subdividing the edges of an original “$3\times 3$ grid”. This edge refinement approach is illustrated by Figure 4 and is inspired by the Multiscale Hybrid-Mixed method [12]. We remark that with both refinement procedures, it is ensured that none of the coarse grids will have a degree of freedom located at the corner $(0,0)$. As a result, the corner singularity will be captured by the basis functions associated with the L-shaped domain. We also note that in order to improve the precision of the fine-scale finite element method, the size of the triangles is graded in the vicinity of the corner $(0,0)$.
Figure 5 reports, for both mesh and edge refinement strategies, the relative error in $H^{1}$ semi-norm and $L^{2}$ norm as functions of maximal coarse edge length $H$. The black dashed line represents the typical fine-scale finite element error; here, we use $\mathbb{P}_{2}$ finite elements. The relative error and the experimental order of convergence for edge refinement procedure are equally reported in Table 1. In accordance with the error estimate (11), for the edge refinement procedure, we observe superconvergence in the energy norm with rates of approximately $3/2$ and $5/2$ for $p=1$ and $p=2$, respectively. In the $L^{2}$ norm, the observed convergence rates are approximately $3$ and $7/2$ for $p=1$ and $p=2$, respectively. In contrast to the edge refinement convergence, the convergence resulting from the mesh refinement seems to be controlled by the low global regularity of the solution. We observe the convergence rates typical for finite element methods on quasi-uniform meshes, that is, close to $2/3$ and $4/3$ in $H^{1}$ and $L^{2}$ norms, respectively.
@@ 4.3 Iterative DD and Preconditioned Krylov Methods on Small Urban Domain
We examine the performance of the two-level stationary iteration and preconditioned GMRES method over a domain based on realistic urban geometries for which the data sets were kindly provided by Métropole Nice Côte d’Azur. We focus here on a relatively small spatial frame shown on the left of Figure 1). The data frame is $160\times 160$ meters and contains two kinds of structural features representing buildings (and assimilated small elevated structures) and walls in urban data. The number of perforations associated to buildings and walls is 63 and 77, respectively. The triangulation of the domain is performed using Triangle [45]. This “small” model domain would take a minimum of 23  000 degrees of freedom to triangulate without imposing maximum triangle area or constraining the mesh to match the coarse partitioning. While the fine-scale mesh will change depending on $N$, the number of degrees of freedom in the mesh will be close to this number. Due to the geometric complexity of the computational domain, the mesh includes a number of very small triangles; the minimal triangle area is given as $5.83\times 10^{-13}$. We consider the model problem (1) with $f=1$, which we discretize with continuous piece-wise affine finite elements; the finite element solution is reported on the left of Figure 1.
@@
We expose on Figure 8 the convergence history of the two-level stationary iteration method with overlap $\frac{1}{20}\mathcal{H}_{j}$. We observe that the convergence of the algebraic error is robust with respect to $h$ and that the fine-scale finite element method (black lines) can be reached in a few steps. Further iterations do not improve the overall precision of the approximate solution even though the algebraic error may decrease. A stopping criterion which prevents excess iterations is warranted; such a posterioiri estimates are discussed for example in [46, 47]. As expected, the performance of the stationary iteration method considerably deteriorates (see Figure 9) in the case of minimal geometric overlap.
@@ 4.4 Scalability Tests for Preconditioned Krylov Method on Larger Urban Domain
For the sake of comparison, we also provide numerical results for the well-known Nicolaides coarse space [48], made of flat-top partition of unity functions associated with the overlapping partitioning. For the Nicolaides coarse space, scalability with respect to the number of subdomains is only guaranteed if the subdomains are connected. Therefore, to generate connected subdomains, we further partition $(\Omega_{j}^{\prime})_{j}$ into a family of connected subdomains. In other words, let $m_{j}$ denote the number of disconnected components for each overlapping subdomain $\Omega_{j}^{\prime}$ and let $\Omega_{j,l}^{\prime},l=1,\ldots,m_{j}$ denote the corresponding disconnected component. Then our new overlapping partitioning contains $m=\sum_{j=1}^{N}m_{j}$ total subdomains and is given by